OpenStudy (anonymous):

what is a function that meets these requirments, no vertical asymptote, horizontal of y=0, y intercept (0,1/4)?

OpenStudy (e.mccormick):

OK, what part of this is giving you trouble?

OpenStudy (anonymous):

the no vertical asymptote part

OpenStudy (anonymous):

No vertical asymptote implies that the domain is all real numbers.

OpenStudy (anonymous):

forgot to mention its a rational function

OpenStudy (e.mccormick):

Rational.... aaah....

OpenStudy (e.mccormick):

Well, then there should be no way for the bottom to be 0.

OpenStudy (anonymous):

x will never be in the denominator.

OpenStudy (e.mccormick):

@EulersEquation What about \[\frac{x^2+\frac{5}{4}x+\frac{1}{4}}{x+1}\]

OpenStudy (tkhunny):

Try \(y = \dfrac{1/4}{x^{2}+ 1}\)

OpenStudy (e.mccormick):

https://www.desmos.com/calculator/oearcetoau

OpenStudy (anonymous):

e.mccormick, x + 1 cannot go in the denominator, otherwise you have a vertical asymptote at x = -1. However, x^2 + 1 is okay. That would be in keeping with a horizontal asymptote at y = 0.

OpenStudy (anonymous):

@e.mccormick, x + 1 cannot go in the denominator, otherwise you have a vertical asymptote at x = -1. However, x^2 + 1 is okay. That would be in keeping with a horizontal asymptote at y = 0.

OpenStudy (e.mccormick):

@EulersEquation Nope. It is a removable. No asymptote.

OpenStudy (e.mccormick):

I was just giving hints. But tkhunny gave the actual solution, so i added it to the graph.

OpenStudy (tkhunny):

@e.mccormick I like it. You could not have done that at x = 0, as this would violate the Intercept requirement. Right, after the bad idea that 'x' could not appear in the denominator, it seemed prudent to get past that.

OpenStudy (anonymous):

\[y = \frac{ 1 }{ x^2 }\] is a good candidate. @tkhunny beat me to it, however.

OpenStudy (anonymous):

@tkhunny gets a medal.

OpenStudy (anonymous):

No -- mine is wrong because that gives a vertical asymptote at x = 0! Dang!

OpenStudy (tkhunny):

That's what the +1 is for. The denominator cannot be zero for Real x.

OpenStudy (anonymous):

Seeing that I am a 17th century mathematician, I am quite old. Getting too old for this sort of thing.

OpenStudy (e.mccormick):

ha! You are not the 17th century mathematician. You are just his brain child! =P

OpenStudy (anonymous):

a wannabe